# Ayoung Lee1, Ryan Sungmo Kwon1, Peter Railton2, Lu Wang1

### v2

@@ Abstract
Navigating high-stakes dilemmas involving conflicting values is challenging even for humans, let alone for AI. Yet prior work in evaluating the reasoning capabilities of large language models (LLMs) in such situations has been limited to everyday scenarios. To close this gap, this work first introduces CLASH  [^1] https://huggingface.co/datasets/launch/CLASH(Character perspective-based LLM Assessments in Situations with High-stakes), a meticulously curated dataset consisting of 345 high-impact dilemmas along with 3,795 individual perspectives of diverse values. In particular, we design CLASH in a way to support the study of critical aspects of value-based decision-making processes which are missing from prior work, including understanding decision ambivalence and psychological discomfort as well as capturing the temporal shifts of values in characters’ perspectives. By benchmarking 10 open and closed frontier models, we uncover several key findings. (1) Even the strongest models, such as GPT-4o and Claude-Sonnet, achieve less than 50% accuracy in identifying situations where the decision should be ambivalent, while they perform significantly better in clear-cut scenarios. (2) While LLMs reasonably predict psychological discomfort as marked by human, they inadequately comprehend perspectives involving value shifts, indicating a need for LLMs to reason over complex values. (3) Our experiments also reveal a significant correlation between LLMs’ value preferences and their steerability towards a given value. (4) Finally, LLMs exhibit greater steerability when engaged in value reasoning from a third-party perspective, compared to a first-person setup, though certain value pairs benefit uniquely from the first-person framing.
@@ 1 Introduction
As large language models (LLMs) become widely used in value-sensitive applications, such as medical (Hu et al., 2024; Singhal et al., 2023; 2025), legal (Nguyen, 2023; Xiao et al., 2021), and finance (Yu et al., 2024; Xie et al., 2023), it is essential that they understand pluralistic values and their nuances (Nagel, 1998; Kekes, 1996; Raz, 1999; James, 1891) to make contextually appropriate judgments. Although progress has been made in recent work (Chiu et al., 2024; Lourie et al., 2021), these studies still focus on everyday situations only. In this paper, we aim to investigate a core research question: Can LLMs make proper judgments in high-stakes dilemmas and from different perspectives? To address this question, we introduce an evaluation framework wherein dilemma situations are presented along with character descriptions representing various perspectives, and LLMs are prompted to make decisions that align with these character-specific perspectives. Within this framework, we assert three major contributions of our work.
@@
Our first contribution is a new dataset, CLASH, for Character perspective-based LLM Assessments in Situations with High-stakes. We define high-stakes conditions as their outcomes carry significant consequences such as loss of life and substantial financial implications. Existing value judgment datasets all focus on everyday dilemmas (Chiu et al., 2024; Lourie et al., 2021) and contain short, up to three sentences situation descriptions (Emelin et al., 2021; Lourie et al., 2021; Hendrycks et al., ). Moreover, many of these datasets rely on synthetically generated situations (Chiu et al., 2024; Scherrer et al., 2024). In contrast, CLASH consists of 345 human-written, long-form, high-stakes dilemmas, with each dilemma being described with about 520 words on average and in high quality. Each dilemma is further enriched with diverse value perspectives, presented as narratives that effectively capture nuances (Fisher, 1984) and context (Li et al., 2024). Our presentation of perspectives more closely reflect how humans engage with LLMs in real-world settings, compared to prior work that only specifies a list of values (Jiang et al., ; Chiu et al., 2024; Sorensen et al., 2024) or characters’ demographics information (Santurkar et al., 2023).
In the second contribution, we leverage the contextualized perspectives in CLASH to examine three important aspects of value reasoning that are previously unexplored: (i) ambivalence for decisions, (ii) psychological discomfort arising from the decision-making process, and (iii) understanding the presence and nature of value shifts. Ambivalence describes the state of indecision caused by competing values (van Delft, 2004). Existing research either constrains LLMs to binary choices in dilemmas (Scherrer et al., 2024; Chiu et al., 2024) or conflates ambivalence with scenario complexity or annotator disagreement (Nie et al., 2023), where the latter can be caused by description ambiguity. On the contrary, we explicitly incorporate an ambivalent setting in CLASH to examine the phenomenon of indecision. Our findings reveal that even the strongest models, such as GPT-4o (Hurst et al., 2024) and Claude-Sonnet (Anthropic, 2024), correctly answer less than half of the cases, highlighting LLMs’ struggle with decision-making ambivalence.
We further investigate psychological discomfort, the internal unease people face in difficult decisions, drawing on cognitive dissonance theory (Festinger, 1957). Recognizing this discomfort is key to assessing whether LLMs understand the emotional weight behind decisions rather than simply producing outcomes. For instance, while a purely utilitarian model might always choose to save the most lives, a more nuanced one would acknowledge the emotional cost of such choices. To our knowledge, this is the first in-depth analysis of psychological discomfort in LLMs’ decision-making process.
Moreover, we study dynamic value shifts as expressed through the character’s perspective, aiming to simulate scenarios where individuals revise their values over time. While prior work like ProgressGym (Qiu et al., 2025) examines long-term societal value shifts, our approach focuses on short-term, individual-level changes within specific narrative contexts. Our results show significant performance drops when LLMs are prompted with evolved values, e.g., up to 66.43% for GPT-4o-mini (Hurst et al., 2024) and 53.62% for Llama3.3-70B (AI, 2024a), highlighting the challenge of reasoning over perspectives with updated values.
@@
Finally, the third contribution of this work is the use of character descriptions in CLASH for evaluating LLMs’ steerability. We focus on how effectively model decisions can be guided by different value pairs, i.e., ‘Can the model follow value X when it conflicts with value Y?’, in contrast to prior work that targets single-value alignment (Dong et al., 2023; Rimsky et al., 2024), such as ‘Can we make the model follow value X?’ This distinction is particularly important for high-stakes dilemmas, where balancing competing values is essential. Our results show a strong negative correlation between model preferences and steerability ($r=–0.329,p<0.0001$). Moreover, framing LLMs as third-party decision-makers generally improves steerability, though some value pairs benefit from first-person framing.
To sum up, we make the following contributions:
  • We curate a set of high-stakes, human-generated, long dilemma situations averaging more than 500 words in high quality, addressing value reasoning in critical yet overlooked scenarios. We also introduce 3,795 character descriptions of different perspectives, encoding context that significantly enrich the complexity of value reasoning in dilemmas.
  • By incorporating diverse contextualized perspectives, we examine previously unexplored aspects in value reasoning: decision ambivalence, psychological discomfort, and value shifts. Results show that LLMs struggle with these nuanced facets, e.g., GPT-4o and Claude-Sonnet achieve below 50% on ambivalent cases, and performance drops by up to 66% when reasoning with value shifts.
  • We examine LLM steerability based on competing value pairs encoded in the character descriptions instead of single values, and uncover a strong negative correlation between model preference and steerability. We also provide practical insights into effective question framing strategies using first-person vs. third-person perspectives.
2 Elements in CLASH
@@
Figure 1: Dataset construction pipeline for CLASH. Key components produced at each step are indicated within the blue boxes. An ”acceptable” value-related rationale means that the person following this rationale would consider the action acceptable. Refer to caption and Refer to caption stand for acceptable rationale and unacceptable rationale, respectively. Refer to Appendix A.3 for a full example.
@@
Figure 2: Categories and the corresponding intended ground-truth answers. The relationship between the two competing value-related rationales within the character description category is intuitively illustrated with the ground-truth answers. Refer to Section 2 for detailed explanations of the categories and questions.
We first introduce four key elements in the benchmark: situation, action, value-related rationales, and character descriptions, with examples presented in Figure 1. Situations are high-stakes dilemmas collected from publicly available web content. Each dilemma involves a hard-to-decide action. We extract value-related rationales supporting both sides of the argument: whether the action is considered acceptable or unacceptable. These rationales are then utilized to create 11 character descriptions in step 3 in Figure 1, spanning 6 categories per dilemma situation. The character’s name is always set to ”Character A” to avoid introducing any bias based on gender, ethnicity, or other personal attributes. We present these character descriptions and ask questions to elicit responses from LLMs, requiring them to select answers based on the character’s perspective.
We first consider character descriptions that do not involve any change in values, which are classified as static and include Straightforward, Simple Contrast, and Swayed Contrast. In contrast, categories that involve varying degrees of value change are assigned as dynamic categories, comprising Shift, Half-Shift, and False-Shift. Each category is described as follows, with intuitive illustrations in Figure 2. In the Straightforward category, one value-related rationale is clearly prioritized over the other, while in the Simple Contrast category, the character equally endorses both rationales. In the Swayed Contrast category, the character acknowledges both rationales but prioritizes one. Characters in the Shift category completely shifts from supporting one rationale to the opposing rationale over time. In the Half-Shift category, the character initially favors one rationale but eventually comes to support both equally. Finally, in the False-Shift category, the character encounters a situation that is likely to challenge their values, but they remain steadfast in their original belief.
@@
For the static categories, we ask two questions: 1) General question: from Character A’s perspective, would the following action—already carried out, being considered to be taken, might be considered to be taken, or being questioned within the context of the situation—be deemed acceptable: {action}? 2) Discomfort question: if the previous answer is yes, would Character A find doing the action mentally or emotionally discomforting? Conversely, if the previous answer is no, would Character A find not doing the action mentally or emotionally discomforting? The first question evaluates the model’s general ability to make a decision based on value perspectives, while the second addresses discomfort triggered by this value-based decision-making process. Refer to Appendix A.1.1 for a detailed discussion on the question design. The general question is a ternary classification task: requiring selection among {”Yes”, ”No”, ”Ambiguous”} [^2] We use the label ”Ambiguous” rather than ”Ambivalent” because ambivalence refers to conflicting feelings experienced directly by a person (a first-person phenomenon) (White, 1978; Zielyk, 1966), whereas our question is framed from a third-person perspective in the main experiments. In this context, ”Ambiguous” reflects the model’s recognition that character holds competing values that make the acceptability of the action indeterminate., while the discomfort question is a binary classification task: requiring selection from {”Yes”, ”No”}.
For dynamic categories, models are presented with two new questions: From Character A’s 1) previous or 2) current value preference, would the following action—already carried out, being considered to be taken, might be considered to be taken, or being questioned within the context of the situation—be deemed acceptable: {action}? These questions are designed to assess the language models’ sensitivity of value shifts. We would refer to the first question as the previous question and the second question as the current question from now on. Both are ternary problems: the model should select from {”Yes”, ”No”, ”Ambiguous”}. Refer to Figure 2 for the ground-truth answers for each category and to Appendix A.1.2 for a detailed explanation of how the ground-truth answers are assigned.
@@ 3 Dataset Construction Procedure for CLASH
Our benchmark curation process consists of four major steps. It is presented in Figure 1 and is described below. Detailed guidelines and corresponding prompts are provided in Appendix A.2.
@@
Step 1) Situation Collection and Action Generation: We first collect high-stakes situations and identify the corresponding hard-to-decide actions. To do so, we identify and crawl websites featuring dilemmas in high-impact domains. The key areas we focus on are: Medical, Business, Journalism/Media, and Government/Politics. We sourced data from five different websites; a comprehensive list is presented in Table 4. The selection criteria for the websites, designed to address potential data contamination and ensure sufficient complexity, are detailed in Appendix A.2.1. Upon compiling the dilemmas, we input these scenarios into GPT-4o to generate the associated hard-to-decide actions relevant to each dilemma. The prompt can be found in Appendix A.2.1. Subsequently, one of the authors manually inspect the results to ensure alignment between the identified dilemmas and generated actions. The full checklist used in this human inspection process is detailed in Appendix A.2.1.
@@
Step 2) Value-Related Rationales Generation: Based on the verified dilemmas and their associated actions, we prompt GPT-4o to generate value-related rationales that either support or oppose the action, with the prompt presented in Appendix A.2.2. Eight human inspectors, who are native English-speaking students majoring in computer science, cognitive science, or linguistics, are then employed to evaluate and refine these rationales, ensuring their relevance to the dilemma and action. The quality of the rationales is maintained through the use of a rigorous checklist, which includes criteria such as clarity, formatting, and relevance to the specific situation. Inspectors revised any rationale that did not meet the checklist criteria. The full checklist is described in Appendix A.2.2.
@@
Step 3) Character Description Generation: We create character descriptions from the value-related rationales. Prompts tailored to each category are used to generate character descriptions via GPT-4o. Refer to Appendix A.2.3 for the specific prompt we used. Five of the human inspectors then refine the character descriptions to ensure the descriptions are aligned with the situations and adhere to the specified category. Refer to Appendix A.2.3 for the full checklist.
@@
Step 4) Dataset Validation: This step involves the annotation of responses to questions, which serves as the validation of our dataset’s quality. The three remaining students who did not participate in the character description inspection engage in the annotation, ensuring they are unaware of the predefined categories and intended ground-truth answers (Figure 2). To validate the inspected character descriptions, an initial annotator reviews the descriptions and provides annotations based solely on their content. This step is crucial for assessing whether the descriptions are clear enough to elicit responses that align with the ground-truth. In cases where the annotation does not align with the intended ground-truth answer, a second annotator independently annotates the answers to the question based on the same character description, functioning as a tie-breaker. Although tie-breaking is typically done by a third person, in our framework the second annotator assumes this role. This is because the inspector of the character description, while not providing an explicit annotation, creates the character description to reflect the intended ground-truth. Thus, the inspector serves as the first implicit annotator providing the intended ground-truth answer as the labels, the first annotator as the second, and the second annotator as the third, effectively acting as the tie-breaker. Excluding tie-breaking cases, each setting is evaluated by exactly one inspector and one annotator. Given a total of five inspectors and three annotators, we compute all pairwise Cohen’s Kappa scores among people and average them. The resulting average score of 0.985 indicates a very high level of inter-annotator agreement, further supporting the reliability of the dataset.
The final dataset consists of 345 distinct situations. Given that each situation included 11 character descriptions as demonstrated in Figure 2, this resulted in a total of 3,795 individual perspectives of diverse values. The distribution across the four topics is as follows: Medical (23.48%), Business (31.88%), Journalism/Media (33.33%), and Government/Politics (11.30%), ensuring a moderate level of coverage across all domains.

### v3

@@ Abstract
Navigating dilemmas involving conflicting values is challenging even for humans in high-stakes domains, let alone for AI, yet prior work has been limited to everyday scenarios. To close this gap, we introduce CLASH (Character perspective-based LLM Assessments in Situations with High-stakes), a meticulously curated dataset consisting of 345 high-impact dilemmas along with 3,795 individual perspectives of diverse values. CLASH enables the study of critical yet underexplored aspects of value-based decision-making processes, including understanding of decision ambivalence and psychological discomfort as well as capturing the temporal shifts of values in the perspectives of characters. By benchmarking 14 non-thinking and thinking models, we uncover several key findings. (1) Even strong proprietary models, such as GPT-5 and Claude-4-Sonnet, struggle with ambivalent decisions, achieving only 24.06 and 51.01 accuracy. (2) Although LLMs reasonably predict psychological discomfort, they do not adequately comprehend perspectives involving value shifts. (3) Cognitive behaviors that are effective in the math-solving and game strategy domains do not transfer to value reasoning. Instead, new failure patterns emerge, including early commitment and overcommitment. (4) The steerability of LLMs towards a given value is significantly correlated with their value preferences. (5) Finally, LLMs exhibit greater steerability when reasoning from a third-party perspective, although certain values (e.g., safety) benefit uniquely from first-person framing.
@@ 1 Introduction
As large language models (LLMs) become widely used in value-sensitive and high-stakes applications such as the ones in medical (Hu et al., 2024; Singhal et al., 2023; 2025), legal (Nguyen, 2023; Xiao et al., 2021), and financial domains (Yu et al., 2024; Xie et al., 2023), it is essential for them to understand pluralistic values and their nuances (Nagel, 1998; Kekes, 1996; Raz, 1999; James, 1891) for contextually appropriate decision making. This paper aims to address a core question: Can LLMs make proper judgments in high-stakes dilemmas according to different perspectives?
@@
Our first contribution is a new dataset, CLASH, which stands for Character perspective-based LLM Assessments in Situations with High-stakes. We define high-stakes conditions as their outcomes carry significant consequences such as loss of life and substantial financial implications. This emphasis marks a fundamental difference from prior work, where they focus on lower-stakes, everyday dilemmas (Chiu et al., 2024; Lourie et al., 2021). CLASH is different from existing value judgment datasets that mostly contain short, up to three sentences situation descriptions (Emelin et al., 2021; Hendrycks et al., ), or rely on synthetically generated situations (Chiu et al., 2024; Scherrer et al., 2024). In contrast, CLASH consists of 345 human-written long-form dilemmas, with each further enriched with diverse value perspectives presented through narratives. Our presentation of perspectives more closely reflects how humans engage with LLMs in real-world settings and effectively captures the nuances inherent in human value systems (Fisher, 1984), compared to prior work that did not consider the contextualization of values (Jiang et al., ; Chiu et al., 2024; Sorensen et al., 2024) or simply do so by including coarse demographic information of characters (Santurkar et al., 2023).
Beyond reaching the right decision, it is crucial, yet often overlooked, to understand the complexities of LLM decision-making for reliable use in high-stakes domains. We address this by leveraging the contextualized perspectives in CLASH to examine three important aspects of value reasoning that are frequently present in real life but previously unexplored: (i) ambivalence between options, (ii) psychological discomfort arising from the decision-making process, and (iii) understanding the presence and nature of value shifts.
Ambivalence describes the state of indecision caused by competing values (van Delft, 2004). Existing research either constrains LLMs to binary choices in dilemmas (Scherrer et al., 2024; Chiu et al., 2024) or conflates ambivalence with scenario complexity or annotator disagreement (Nie et al., 2023). In contrast, we explicitly incorporate ambivalent settings in CLASH to examine the phenomenon of indecision. Our findings reveal that even the strongest proprietary thinking models such as GPT-5 (OpenAI, 2025) and Claude-4-Sonnet (Anthropic, 2025) achieve only 24.06 and 51.01 accuracy, highlighting the struggle of LLMs with ambivalent decisions. Moreover, we are the first to investigate psychological discomfort, the internal unease people face in difficult decisions, drawing on cognitive dissonance theory (Festinger, 1957). Next, we study dynamic value shifts as expressed through the perspective of characters, aiming to simulate scenarios where individuals revise their values over time. Our results show significant performance drops when LLMs are prompted with shifted values, with an average accuracy drop of 42.95 points and a maximum of 66.43 points for GPT-4o-mini (Hurst et al., 2024), highlighting the challenge of reasoning over perspectives with updated values.
Beyond measuring accuracy, we identify features of successful and unsuccessful reasoning chains in thinking models. Backchaining and verification, effective in math and game reasoning domains (Gandhi et al., 2025), prove less effective in value-based reasoning, while new patterns including early commitment and overcommitment frequently characterize failures. We also analyze the chains through ethical theories, suggesting that pragmatic ethics (Dewey & Tufts, 2022) may serve as a promising framework for eliciting more effective reasoning.
@@
Finally, we introduce a method for measuring a newly proposed concept: conditional steerability that evaluates how effectively a model can be guided toward one value when it conflicts with another, which captures the true tensions of real-world dilemmas. This stands in contrast to prior work on absolute steerability, which focuses on single-value alignment (Dong et al., 2023; Rimsky et al., 2024). Our results show a significant negative correlation between model preferences and steerability (r=0.243,p<0.005r=-0.243,p<0.005). Furthermore, we use conditional steerability to examine the effect of question framing. We find that framing LLMs as third-party decision-makers generally improves steering, though certain value pairs benefit more from first-person framing.
2 Design of CLASH
@@
Figure 1: Dataset construction pipeline for CLASH. Key components produced at each step are indicated within the blue boxes. Refer to caption and Refer to caption stand for supporting rationale and opposing rationale, respectively. Refer to Appendix A.3 for a full example.
@@
Figure 2: Categories and the corresponding intended ground-truth answers. The relationship between the supporting and opposing value-related rationales within the character description category is intuitively illustrated with the ground-truth answers.
Elements in CLASH.
CLASH consists of four key components: 1) situation, 2) action, 3) value-related rationales, and 4) character descriptions. Figure 1 depicts the overall structure and generation pipeline of CLASH. Situations are high-stakes dilemmas curated from public web sources, which include rich details and reflect real-life complexities in decision making. For example, a kidney transplant decision may become more difficult if the doctor’s close friend is one of the patients, or if time pressure requires choosing quickly to ensure at least one life can be saved. We further analyze the impact of these contextual details in Appendix A.2, where we show that the performance differences between the detailed and simplified settings are statistically significant. This finding demonstrates that our proposed setting constitutes a distinct task relative to prior approaches (Emelin et al., 2021; Hendrycks et al., ) that presents simple narratives in no more than three sentences. Each dilemma involves a corresponding controversial action, and we extract value-related rationales that either support or oppose the action. These rationales are then utilized to create character descriptions. To avoid bias from gender, ethnicity, or other personal attributes, the character is always named as “Character A.” We present these descriptions to LLMs and ask questions that require answers from the character’s perspective.
Character Description Categories.
Character descriptions are classified into two large categories (static and dynamic), each with three subcategories (Figure 2). Character descriptions that exhibit no change in values are classified as static and include Straightforward, Simple Contrast, and Swayed Contrast. In contrast, character descriptions that entail varying degrees of value change are assigned as dynamic categories, comprising Shift, Half-Shift, and False-Shift. Each subcategory has two descriptions except for Simple Contrast, resulting in a total of 11 character descriptions per situation.
@@
For the static categories, we pose two questions: 1) General question: “From Character A’s perspective, would the following action—already carried out, being considered to be taken, might be considered to be taken, or being questioned within the context of the situation—be deemed acceptable: {action}?” 2) Discomfort question: “If the previous answer is yes, would Character A find doing the action mentally or emotionally discomforting? Conversely, if the previous answer is no, would Character A find not doing the action mentally or emotionally discomforting?” The first question evaluates the general ability of the model to make a decision based on value perspectives, while the second addresses discomfort triggered by this value-based decision-making process. Refer to Appendix A.1 for a detailed discussion on the question design. The general question is a ternary classification task, requiring selection among {“Yes”, “No”, “Ambiguous”} [^1] We use the label “Ambiguous” rather than “Ambivalent” because ambivalence refers to conflicting feelings experienced directly by a person (a first-person phenomenon) (White, 1978; Zielyk, 1966), whereas our question is framed from a third-person perspective in the main experiments. In this context, “Ambiguous” reflects the model’s recognition that character holds competing values that make the acceptability of the action indeterminate., while the discomfort question is a binary task, requiring selection from {“Yes”, “No”}. For dynamic categories, models are presented with two new questions: “From Character A’s 1) previous or 2) current value preference, would the following action be deemed acceptable: {action}?” The questions are designed to assess the sensitivity of language models to value shifts, and both are ternary problems.
For the static categories, Straightforward involves one rationale clearly dominating the other, eliciting a clear-cut general answer without discomfort. In Simple Contrast, both rationales are equally endorsed, resulting in an “Ambiguous” response with no evaluation of the discomfort question. The Swayed Contrast represents a middle ground; while both rationales are acknowledged, one is prioritized, resulting in an unambiguous “Yes” or “No” for the general question but with discomfort. For the dynamic categories, Shift shows a complete change of preference within rationales, resulting in opposite answers for previous and current questions. In Half-Shift, an initially favored rationale gives way to equal endorsement, thereby yielding an “Ambiguous” current response. Finally, False-Shift characters are exposed to potentially value-challenging context but remain steadfast in their original belief, producing the same answers for both questions. The intuitive illustration is presented in Figure 2 and an example for each category is provided in Figure 7 in Appendix A.3.
@@ 3 Dataset Construction Procedure for CLASH
The benchmark curation process comprises four steps, presented in Figure 1 and detailed below.
@@
Step 1) Situation Collection and Action Generation. We first collect situations by identifying and crawling websites featuring dilemmas in high-impact domains; a comprehensive list is presented in Table 3. The scenarios are inputted into GPT-4o to generate the associated action relevant to each dilemma. Subsequently, one of the authors manually inspects the results to ensure alignment between the dilemmas and the generated actions. The selection criteria for the websites, the prompt used to generate the actions, and the full checklist for human inspection are detailed in Appendix B.1.
@@
Step 2) Value-Related Rationales Generation. Based on verified dilemmas and their associated actions, we prompt GPT-4o to generate value-related rationales that either support or oppose the action. Eight native English-speaking students are then employed to evaluate and refine these rationales, ensuring their relevance to the dilemma and action. The quality of the rationales is maintained through the use of a rigorous checklist, which includes criteria such as clarity, formatting, and relevance to the specific situation. The inspectors revise any rationale that did not meet the checklist criteria. The prompt and checklist used in this section are presented in Appendix B.2.
@@
Step 3) Character Description Generation. We generate character descriptions from value-related rationales using prompts tailored to each category. These initial descriptions are produced with GPT-4o, then reviewed and refined by five inspectors to ensure they align with the given situations and conform to the specified category. For details on the exact prompt and checklist, see Appendix B.3.
@@
Step 4) Dataset Validation. This step involves the annotation of the responses to questions, which serves as the validation of the quality of CLASH. Three students who are unaware of the predefined categories and intended ground-truth answers engage in the annotation. This step is crucial for assessing whether the descriptions are clear enough to elicit responses that align with the ground-truth. We compute all pairwise Cohen’s Kappa scores among people and average them. The resulting average score of 0.985 indicates a very high level of inter-annotator agreement, further supporting the reliability of the dataset. The detailed process is explained in Appendix B.4.
The final dataset consists of 345 distinct situations. Given that each situation included 11 character descriptions as demonstrated in Figure 2, this results in a total of 3,795 individual perspectives of diverse values. We further recruit two annotators who grew up outside the United States to indicate whether the events described in each dilemma are specific to US. The annotations show that only 21% of the events were US-focused, with a moderate Cohen’s Kappa score of 0.458. This suggests that, despite being in English, CLASH is not exclusively centered on US culture.